Hoi Chong Transport, Ltd. operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is driven 146,000 kilometres during a year, the average operating cost is 12.9 cents per kilometre. If a truck is driven only 73,000 kilometres during a year, the average operating cost increases to 14.8 cents per kilometre. (The Singapore dollar is the currency used in Singapore.)

Required:
1.

Using the high-low method, estimate the variable and fixed cost elements of the annual cost of truck operation. (Round the "Variable cost" to 3 decimal places.)

variable cost = per kilometer

fixed cost= per year

  

2.

Express the variable and fixed costs in the form Y = a + bX. (Round the "Variable cost" to 3 decimal places.)

y= + x

      

3.

If a truck were driven 83,000 kilometres during a year, what total cost would you expect to be incurred? (Round your intermediate calculations to 3 decimal places.)

total annual cost=
Answer #1
1 Kilometre driven Average operating cost (in cents) Total operating cost
a b a*b
146000 0.129 18834
73000 0.148 10804
Variable cost=Change in total operating cost/Change in kilometres driven=(18834-10804)/(146000-73000)=8030/73000=0.11=11 cents per Kilometre
Take 146000 kilometres driven
Fixed cost=Total operating cost-Variable cost=18834-(146000*0.11)=18834-16060=$ 2774 per year
2 Y=2774+0.11X
3 Total annual cost=Fixed cost+(Kilometres driven*Variable cost)=2774+(83000*0.11)=2774+9130=$ 11904